The mean weight of luggage checked by a randomly selected economic class passenger flying between two cities on a certain airline is 40Ib, and the standard deviation is 10 lb. The mean and standard deviation for a business class passenger are 30lb and 6 lb respectively. If there are 50 economic class and 12 business class passengers on the flight, what is the variance of total luggage weight?

Select one:  a. 2360  b. 5000  c. 432  d. 5432 


1
Expert's answer
2021-10-29T03:00:57-0400

The mean luggage weight of economic class passengers is 40 LB and the standard deviation is 10 LB.

The mean luggage weight of business class is 30 LB and the standard deviation is 6 LB.

Therefore, the mean of 12 business class and 50 economic class passengers is "(50 \times 40+12 \times 30)=560"

"variance = 50 \times 10^2+12 \times 6^2=5432"

Answer: d. 5432